### v2

# k-LLMmeans: Scalable, Stable, and Interpretable Text Clustering via LLM-based Centroids

Abstract
We introduce k-LLMmeans, a novel modification of the k-means algorithm for text clustering that leverages LLM-generated summaries as cluster centroids, capturing semantic nuances often missed by purely numerical averages. This design preserves the core optimization properties of k-means while enhancing semantic interpretability and avoiding the scalability and instability issues typical of modern LLM-based clustering. Unlike existing methods, our approach does not increase LLM usage with dataset size and produces transparent intermediate outputs. We further extend it with a mini-batch variant for efficient, real-time clustering of streaming text. Extensive experiments across multiple datasets, embeddings, and LLMs show that k-LLMmeans consistently outperforms k-means and other traditional baselines and achieves results comparable to state-of-the-art LLM-based clustering—with a fraction of the LLM calls. Finally, we present a case study on sequential text streams and introduce a new benchmark dataset constructed from StackExchange to evaluate text-stream clustering methods.
1 Introduction
Text clustering is a fundamental task in natural language processing (NLP), widely applied in document organization, topic modeling, and information retrieval [37, 42]. A common approach involves generating text embeddings [13, 36, 27, 32, 8, 21] for each document, which are then clustered using traditional algorithms [33]. Among these, k-means [26] is one of the most widely used, iteratively refining cluster centroids based on the mean of assigned points. Several alternative methods exist for computing centroids [19, 7, 23], yet they often rely on purely numerical calculations. While effective, such approaches risk losing crucial contextual and semantic information, as averaging high-dimensional embeddings can obscure nuanced textual relationships [34].
We propose k-LLMmeans, a modification to the k-means algorithm by leveraging large language models (LLMs) to dynamically generate cluster centroids. Instead of always computing centroids as the mean of embedding vectors (k-means), in spaced iterations we generate centroids based on textual summaries of the cluster contents. At each of these iterations, the centroid is represented as the embedding of the LLM-generated summary of the texts belonging to the cluster (summarization step). LLM-based centroids preserve key contextual and semantic aspects of the main documents in the cluster while filtering out secondary or lower-priority documents [20]. Figure 1 illustrates the k-LLMmeans algorithm with a single summarization step. The example demonstrates how this step enables the algorithm to escape local minima (common in k-means) and converge to the optimal solution.
Unlike k-means, k-LLMmeans generates a concise textual summary for every centroid at each summarization step, exposing the semantic drift of clusters in real time. This built-in transparency simplifies debugging and validation and is especially valuable in sequential or streaming scenarios, where clusters evolve over time. Figure 2 illustrates one centroid through five summarization updates. To capitalize on this property for large-scale or online data, we adapt mini-batch k-means [38] by inserting our summarization step into a mini-batch k-LLMmeans.
Several studies have explored unsupervised text clustering using LLMs [51, 15, 10, 45, 39, 43, 28], demonstrating state-of-the-art performance across various datasets and benchmarks. These approaches consistently outperform traditional clustering algorithms. Yet despite their impressive empirical gains, existing LLM-based clustering pipelines still face two practical obstacles that limit widespread adoption. First, lack of scalability. Most methods rely on fine-tuning [51], semi-supervised approaches [45], or extensive LLM queries that grow at least linearly with sample size [15, 10], quickly becoming prohibitive for real-world applications. Second, unclear optimization landscape. Most methods braid together bespoke heuristics—hard prompts, similarity thresholds, or greedy merges—without a single, well-defined loss function, leaving convergence behaviour opaque and theoretical guarantees out of reach. We address these two shortcomings with k-LLMmeans: First, by summarizing each cluster in batches we cap LLM usage independent of dataset size; and second, by inheriting the classic k-means objective we retain a provable descent property and a transparent optimization landscape. Regardless of summarization quality, the algorithm will converge to a local minimum.
Our approach is not intended to outperform complex state-of-the-art LLM-based clustering methods, but rather to provide a robust, scalable and LLM-efficient enhancement to the well-established k-means algorithm, preserving its core behavior while using significantly fewer LLM calls than contemporary LLM-based clustering approaches. Fundamentally, a centroid is a numerical abstraction that represents a cluster, and we posit that an LLM-generated textual summary can serve an analogous role. Crucially, our approach preserves most of the mathematical properties of k-means, ensuring that core theoretical guarantees—such as convergence behavior, cluster compactness, and complexity— remain intact.
Figure 1: Illustration of a single-summarization k-LLMmeans algorithm. The first panel displays the text embeddings projected onto the first two principal components along with sample texts, with stars indicating initial random centroids. The second panel shows the local minimum reached by standard k-means. The third panel illustrates the summarization step, where centroids are updated based on LLM-generated summaries of the clustered texts. The final panel presents the resulting clusters after the final k-means step, showing convergence to a more optimal k-LLMmeans solution.
In summary our contributions are as follows:
  • In Section 3 we introduce k-LLMmeans, which enhances the classical k-means algorithm with LLM-powered centroid updates for robust text clustering; Section 4 then builds on this foundation with a mini-batch variant that brings sequential, scalable, and intrinsically interpretable clustering to streaming text data.
  • Section 5 presents extensive simulations demonstrating that k-LLMmeans consistently surpasses k-means and other traditional clustering baselines across a wide range of embeddings and LLMs, while obtaining comparable results to recent LLM-based clustering methods with substantially fewer LLM calls.
  • Section 6 shows that mini-batch k-LLMmeans decisively outperforms streaming variants of k-means, and the case study in Section 7 illustrates how its evolving centroids provide clear, human-readable insight into real-time text streams.
As a by-product of our experiments, we compile a dataset from StackExchange [41] suitable for benchmarking text-streaming analysis methods (Appendix A.2).
Figure 2: An illustration of the dynamic evolution of an LLM-generated centroid during the k-LLMmeans clustering process.
2 Preliminaries: k-means for text clustering
We can formalizae the text clustering problem. Given a corpus of $n$ text documents $D=\{d_{1},\cdots d_{n}\}$. Each document $d_{i}$ is represented as a $d$-dimensional embedding vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$ such that:
$\mathbf{x}_{i}=\text{Embedding}(d_{i}).$
The goal of k-means clustering is to partition these $n$ document embeddings into $k$ clusters, minimizing the intra-cluster variance. Formally, we define the clustering objective as:
$\min_{C_{1},C_{2},\dots,C_{k}}\sum_{j=1}^{k}\sum_{i\in[C_{j}]}\|\mathbf{x}_{i}-\boldsymbol{\mu}_{j}\|^{2},$ (1)
where $C_{j}$ denotes the set of embeddings assigned to cluster $j$, $[C_{j}]=\{i|\mathbf{x}_{i}\in C_{j}\}$ denotes the set of embedding indices assigned to cluster $j$ and $\boldsymbol{\mu}_{j}$ is the cluster centroid, computed as the mean of the assigned embeddings:
$\boldsymbol{\mu}_{j}=\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}.$ (2)
The k-means algorithm assigns each document embedding $\mathbf{x}_{i}$ to the closest centroid based on the smallest distance and updates centroids accordingly until convergence after $T$ iterations. The primary objective of the k-means algorithm is to iteratively adjust cluster centroids to minimize the within-cluster variance, effectively guiding the algorithm toward an optimal set of centroids. However, due to its sensitivity to initialization and the non-convex nature of its objective function, k-means does not guarantee convergence to the global optimum and can instead become trapped in local optima [26, 25]. Various strategies, such as k-means++ initialization and multiple restarts, have been proposed to mitigate these issues and improve the likelihood of achieving better clustering results [4].
3 k-LLMmeans
input : $\ D=\{d_{1},\dots,d_{n}\},k,I,m,l,T$
for $i\leftarrow 1$ to $n$ do
       $\mathbf{x}_{i}=\text{Embedding}(d_{i})$;
end for
for $t\leftarrow 1$ to $T$ do
       if $t=1$ then
             // Initialize using k-means++ (omitted if already initialized)
             $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\}\leftarrow\texttt{k-means++}(\{d_{1},\dots,d_{n}\},k)$;
       end if
      else if $t\bmod l=0$ then
             // Update with LLM every $l$ iterations (summarization step)
             for $j\leftarrow 1$ to $k$ do
                   $m_{j}\leftarrow\min(m,|C_{j}|)$;
                  $\{d_{z_{1}},\dots,d_{z_{m_{j}}}\}\leftarrow\texttt{k-means++}(\{d_{i}\mid i\in[C_{j}]\},m_{j})$;
                  $p_{j}\leftarrow\text{Prompt}(I,\{d_{z_{1}},\dots,d_{z_{m_{j}}}\})$;
                   $s_{j}\leftarrow f_{\text{LLM}}(p_{j})$;
                   $\boldsymbol{\mu}_{j}\leftarrow\text{Embedding}(s_{j})$;
             end for
       end if
      else
             // Update centroids using standard averaging (k-means step)
             for $j\leftarrow 1$ to $k$ do
                   $\boldsymbol{\mu}_{j}\leftarrow\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}$;
             end for
       end if
      for $j\leftarrow 1$ to $k$ do
             $C_{j}=\{\}$;
       end for
      for $i\leftarrow 1$ to $n$ do
             $j^{*}\leftarrow\arg\min_{j\in\{1,\dots,k\}}d(x_{i},\boldsymbol{\mu}_{j})$;
             // Assign $x_{i}$ to cluster $C_{j^{*}}$
             $C_{j^{*}}\leftarrow C_{j^{*}}\cup\{x_{i}\}$;
       end for
end for
return $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\},\{s_{1},\dots,s_{k}\}$
Algorithm 1 k-LLMmeans
To enhance k-means for text clustering, we introduce k-LLMmeans, a novel variant that integrates LLM-based centroids at intermediate stages of the clustering process. The formal procedure is outlined in Algorithm 1. The key distinction between k-LLMmeans and standard k-means lies in the centroid update mechanism: every $l$ iterations, the traditional update step in Equation (2) is replaced by
$\boldsymbol{\mu}_{j}=\text{Embedding}(f_{\text{LLM}}(p_{j}))\$ (3)
where $p_{j}=\text{Prompt}\left(I,\{d_{z_{i}}|z_{i}\sim[C_{j}]\}_{i=1}^{m_{j}}\right)$, and $m_{j}=\min(m,|C_{j}|)$. Here, $z_{i}\sim[C_{j}]$ denotes a sampled index of the embeddings assigned to cluster $C_{j}$ (without repetitions) and $m$ is a parameter that represents the maximum number of sampled indices used to compute the cluster centroid $\boldsymbol{\mu}_{j}$. In simple terms, we update a cluster’s centroid by using the embedding of the response generated by an LLM when queried with a prompt containing a summarization instruction $I$ and a representative sample of documents from the cluster. Rather than providing all documents within the cluster as input, the LLM processes a representative sample as a context prompt. While incorporating the entire cluster is theoretically possible, it poses practical challenges due to prompt length limitations. Therefore, we propose selecting the sample cluster documents using a k-means++ sampling of the cluster embeddings. Our experiments demonstrate that this sampling process facilitates a more effective synthesis of the cluster’s content, leading to improved summaries and, consequently, more refined centroid updates. The instruction $I$ varies depending on the clustering task, but standard summarization prompts are generally sufficient. Figure 1 illustrates how k-LLMmeans augments the standard k-means algorithm by introducing few summarization steps between clustering iterations.
Advantages over k-means. This novel approach can help mitigate the tendency of k-means to get stuck in local optima by dynamically adjusting centroids using semantic insights from the data. Unlike standard k-means, which relies solely on Euclidean updates, an LLM can refine centroids based on contextual meaning and high-dimensional representations, allowing for better adaptation to complex structures in text data. By periodically re-centering clusters with LLM-informed embeddings, the algorithm can escape poor local minima and achieve more coherent and semantically meaningful clusters, even when the initial k-means++ seeding is suboptimal. Apart from this modification, our algorithm adheres to the core principles of k-means, preserving its well-established properties and ensuring practical robustness. We demonstrate in Section 5.1 that k-LLMmeans consistently outperforms k-means across extensive simulations.
Advantages over advanced LLM-based clustering methods. Our approach offers three key advantages over more complex LLM-based clustering methods: (1) Optimization landscape. Grounded in the standard k-means objective, k-LLMmeans inherits provable convergence without relying on the fragile heuristics common in newer LLM-driven methods. A faulty summary merely returns the algorithm to a regular k-means local optimum, whereas competing approaches hinge on stable, format-specific LLM outputs. (2) Scalability. Unlike most state-of-the-art methods, whose LLM usage complexity grows with sample size [15, 10] or require fine-tuning [51], our algorithm only performs $O(kl)$ LLM calls, with even small $l$ values yielding substancial performance gains. (See Section 5.1). (3) Interpretability. By replacing numerical centroids with LLM-generated textual summaries, k-LLMmeans turns each centroid into a concise, human-readable synopsis of its cluster. Practitioners can therefore monitor how cluster semantics evolve across iterations without any post-hoc labeling or visualization. While comparable insights can be extracted indirectly from other clustering algorithms, k-LLMmeans embeds this interpretability natively. Figure 2 illustrates how these summaries progressively sharpen as the algorithm converges.
4 Mini-batch k-LLMmeans
input : $\ \{D_{1},\cdots,D_{b}\},k,I,m,l,T$ // $b$ batches of documents
for $j\leftarrow 1$ to $k$ do
       $C_{j}=\{\}$;
end for
$\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\}\leftarrow\{{\bf 0},\dots,{\bf 0}\}$;
for $i\leftarrow 1$ to $b$ do
// Compute k-LLMmeans with documents in batch11      1Here k-LLMmeans is initialized with the final centroids of the previous batch
       $\{\boldsymbol{\mu}^{*}_{1},\dots,\boldsymbol{\mu}^{*}_{k}\},\{C^{*}_{1},\dots,C^{*}_{k}\},S_{b}\leftarrow\texttt{k-LLMmeans}(D_{i},k,I,m,l,T)$;
       // Update centroids proportional to cluster and batch sizes
       for $j\leftarrow 1$ to $k$ do
             $\eta\leftarrow\frac{|C^{*}_{j}|}{|C_{j}|+|C^{*}_{j}|}$;
             $\boldsymbol{\mu}_{j}\leftarrow\boldsymbol{\mu}_{j}(1-\eta)+\eta\boldsymbol{\mu}^{*}_{j}$;
       end for
end for
return $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\},\{S_{1},\dots,S_{b}\}$
Algorithm 2 Mini-batch k-LLMmeans
Mini-batch k-means [38] is an efficient strategy for large-scale text clustering that processes small, randomly sampled mini-batches instead of the full dataset. This approach substantially reduces memory usage and computational cost, making it well suited for continuously generated text streams—such as those from social media, news, or customer feedback—where data must be clustered incrementally without full dataset access. Mini-batch k-means exhibits convergence properties comparable to standard k-means while offering superior scalability.
Although numerous streaming clustering methods that do not rely on LLMs have been studied [40, 2, 35, 3, 1, 31], only a few have incorporated LLMs [43, 28]. Moreover, existing offline LLM-based clustering approaches face scalability issues, highlighting the need for scalable LLM-driven clustering in an online setting. To address this gap, we introduce mini-batch k-LLMmeans, which directly extends mini-batch k-means by incorporating minimal LLM usage during centroid updates. Algorithm 2 details how mini-batch k-LLMmeans sequentially receives $b$ batches of documents $D_{1},\dots D_{b}$ where each batch contains a set of documents (these batches can either be random samples from a large corpus or represent sequential data). It processes each batch sequentially with k-LLMmeans and updates centroids incrementally using a weighted rule like mini-batch k-means. Our mini-batch k-LLMmeans algorithm preserves the desirable properties of mini-batch k-means, with low memory and LLM usage. Section  6 shows that it also outperforms in simulations.
5 Static experiments
We use four benchmark datasets that span diverse domains and classification granularities: Bank77 [9], CLINC [24], GoEmo [12] and MASSIVE (domain and intent) [16]. See Appendix A.1 for detailed descriptions of the datasets. We evaluate our k-LLMmeans algorithm on each of the four static (non-streaming) dataset using the known number of clusters and performing 120 centroid-update iterations. To demonstrate the robustness of our approach, we compute embeddings with four different pretrained models: DistilBERT [36], e5-large [46], S-BERT [34], and text-embedding-3-small [30]. For the LLM component, we use GPT-3.5-turbo, GPT-4o, Llama-3.3, Claude-3.7 and DeepSeek-V3.
For the instruction task $I$, we employ a simple summarization prompt that adapted to each dataset. For example, for Bank77 we use the prompt: “The following is a cluster of online banking questions. Write a single question that represents the cluster concisely.” We evaluate four variants of k-LLMmeans, differing in the number of summarization steps and the prompt configuration. The single variant uses a single summarization step with all cluster documents as input ($l=60$), while the multiple variant performs five summarization steps with shorter prompts ($l=20$), again using all cluster documents each time. Additionally, we explore few-shot variants (FS-single and FS-multiple), where each prompt includes $m=10$ randomly selected documents instead of the full cluster.
As traditional clustering baselines, we consider k-means, k-medoids [23], spectral clustering [29], agglomerative clustering [22], and Gaussian Mixture Models (GMM) [11]. We use ground-truth number of clusters and ten different seeds. For advanced LLM-based methods, we compare with results obtained by Feng et al. [15] on ClusterLLM [51], IDAS [10], and LLMEdgeRefine [15].
5.1 Results with static data
Table 1: Average Accuracy (ACC) and Normalized Mutual Information (NMI) for four k-LLMmeans variants using GPT-4o, compared against traditional clustering baselines, with text-embedding-3-small embeddings, evaluated on four benchmark datasets.
Dataset/Method BANK77 CLINC GoEmo Massive (D) Massive (I)
ACC NMI ACC NMI ACC NMI ACC NMI ACC NMI
k-LLMmeans
   single 67.1 83.6 78.1 92.5 24.0 22.3 61.1 69.6 54.0 73.0
   multiple 66.7 83.6 79.1 92.8 24.1 22.1 60.6 69.5 55.8 73.5
   FS-single 67.5 83.8 79.2 92.8 23.0 21.9 60.8 69.4 55.3 73.6
   FS-multiple 67.9 84.1 80.2 93.1 24.2 22.3 63.2 70.6 56.4 73.9
k-means 66.2 83.0 77.3 92.0 20.7 20.5 59.4 67.9 52.9 72.4
k-medoids 41.7 69.5 49.3 77.7 15.7 15.9 37.6 38.8 35.9 52.4
GMM 67.7 83.0 78.9 92.7 21.5 20.6 56.2 68.6 53.9 73.2
Agglomerative 69.9 83.7 81.0 92.5 15.8 14.1 62.8 67.1 56.6 70.5
Spectral 68.2 83.3 76.3 90.9 17.6 15.2 61.5 67.0 56.7 71.4
Table 2: Average ACC, NMI, and distance between final and true centroids ($dist$) for k-means and k-LLMmeans FS-multiple, evaluated on three datasets using four different embedding models.
Dataset CLINC GoEmo Massive (D)
embedding/Method ACC NMI $dist$ ACC NMI $dist$ ACC NMI $dist$
DistilBERT
   k-means 53.9 77.2 0.340 17.8 18.3 0.364 44.4 45.1 0.309
   k-LLMmeans 55.3 78.7 0.343 18.2 18.8 0.351 46.2 46.4 0.295
text-embedding-3-small
   k-means 77.3 92.0 0.200 20.7 20.5 0.287 59.4 67.9 0.246
   k–LLMmeans 80.2 93.1 0.179 24.2 22.3 0.278 63.2 70.6 0.227
e5-large
   k-means 73.8 90.8 0.131 22.8 22.8 0.176 58.4 63.7 0.138
   k-LLMmeans 77.2 92.5 0.119 24.2 24.3 0.168 62.3 65.9 0.133
S-BERT
   k-means 76.9 91.0 0.215 13.7 13.3 0.355 58.2 64.6 0.271
   k-LLMmeans 79.7 92.5 0.198 14.7 13.9 0.346 59.7 65.6 0.254
Table 3: Number of LLM calls (prompts), average ACC, and average NMI for k-LLMmeans FS-multiple using various LLMs and e5-large embeddings, compared against other LLM-based text clustering methods on three benchmark datasets.
Dataset/Method CLINC GoEmo Massive (D)
prompts ACC NMI prompts ACC NMI prompts ACC NMI
k-LLMmeans
   GPT-3.5 750 76.0 92.1 135 23.9 24.1 90 63.3 66.2
   GPT-4o 750 77.2 92.5 135 24.2 24.3 90 62.3 65.9
   Llama-3.3 750 77.2 92.3 135 23.8 23.1 90 62.0 66.3
   DeepSeek-V3 750 69.7 90.8 135 22.8 22.7 90 62.6 66.3
   Claude-3.7 750 76.9 92.5 135 24.2 23.7 90 61.8 66.0
ClusterLLM 1618 83.8 94.0 1618 26.8 23.9 1618 60.9 68.8
IDAS 4650 81.4 92.4 3011 30.6 25.6 2992 53.5 63.9
LLMEdgeRefine 1350 86.8 94.9 895 34.8 29.7 892 63.1 68.7
Comparing k-LLMmeans variants and traditional clustering methods. Table 1 reports Average Accuracy (ACC) and Normalized Mutual Information (NMI) for four k-LLMmeans variants using GPT-4o, compared against traditional clustering baselines, with text-embedding-3-small embeddings, evaluated on four benchmark datasets. Overall, all k-LLMmeans variants outperform traditional clustering methods in terms of NMI and in general achieve higher average accuracy across datasets. Within k-LLMmeans variants there are minor differences, although running multiple summarization steps tends to offer better performance. Interestingly, using few-shot summarization generally leads to better results than prompting on the entire cluster. These observations suggest k-LLMmeans improves over k-means while remaining efficient, as few-shot summarization appears sufficient without needing extensive prompt length (See Appendix B for standard errors of all simulations).
[…]
Comparing k-LLMmeans with different LLMs and state-of-the-art LLM-based clustering methods. Table 2 reports number of LLM calls (prompts) and Average Accuracy (ACC) and Normalized Mutual Information (NMI) for the k-LLMmeans FS-multiple variant on three benchmark datasets, pairing e5-large embeddings with several backbone LLMs. Performance is stable across LLM choices, indicating that k-LLMmeans is robust to the specific language model. For reference, Table 2 also reports results from Feng et al. [15] for ClusterLLM, IDAS, and LLMEdgeRefine. All three baselines rely on GPT-3.5: ClusterLLM (with fine-tuning) and IDAS use the same e5-large embeddings as we do, whereas LLMEdgeRefine benefits from the stronger Instructor embeddings [21]. Even so, our GPT-3.5 configuration attains ACC and NMI that are only marginally lower—and often comparable—while requiring orders of magnitude fewer LLM calls and no fine-tuning. This substantial reduction in model prompts translates directly into improved scalability and cost-efficiency.
6 Sequential (mini-batch) experiments
We use yearly posts from 2020 to 2023 from 35 StackExchange sites [41]. Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods (See Appendix A.2 for details). For each yearly subset, we split the data into $b=\lceil\tfrac{n}{10000}\rceil$ equal-sized batches ${D_{1},\dots,D_{b}}$ in chronological order, where $n$ is the number of documents for that year. We then use the ground-truth clusters and run the mini-batch k-LLMmeans algorithm in its four variants described in Section 5 (for practical reasons for the full cluster method we use 50 samples). For comparison, we evaluate three baselines: mini-batch k-means with standard random sampling across the entire year (until convergence), sequential mini-batch k-means with $b$ chronological batches, and standard k-means on the full dataset. In all experiments, we use text-embedding-3-small and GPT-4o for the summarization step.
6.1 Results with streaming data
Table 4: Average ACC, and average NMI for four sequential mini-batch variants, k-means, mini-batch k-means, sequential mini-batch k-means on the yearly StackExchange data.
Year/Method 2020 2021 2022 2023
(69147 posts) (54322 posts) (43521 posts) (38953 posts)
ACC NMI ACC NMI ACC NMI ACC NMI
mini-batch k-LLMmeans
   single 73.1 81.2 73.2 80.3 70.8 79.4 72.3 80.0
   multiple 72.5 80.9 73.6 80.5 74.4 80.5 73.4 80.3
   FS + single 74.0 81.1 72.2 79.8 70.2 79.3 71.6 79.8
   FS + multiple 75.4 81.6 73.5 80.2 72.8 80.1 72.7 80.1
k-means 73.4 80.6 67.7 79.0 68.6 79.0 72.0 79.6
mini-batch k-means 67.0 78.2 67.7 77.4 67.5 77.6 67.2 77.0
seq. mini-batch k-means 67.0 76.6 66.7 75.2 65.6 75.6 65.8 74.8
Table 4 presents ACC and NMI on the yearly StackExchange corpus, comparing four sequential mini-batch variants of k-LLMmeans against standard k-means, mini-batch k-means, and sequential mini-batch k-means across five random seeds. The mini-batch k-LLMmeans methods consistently outperform all three baselines across all scenarios. Notably, they even surpass standard k-means, which operates on the full dataset rather than mini-batches. This highlights the scalability and efficiency of our approach for sequential data, achieving superior performance despite processing data sequentially. As observed previously, few-shot summarization proves effective, mitigating the need for extensive LLM usage.
7 Case study
To demonstrate the interpretability of our method in capturing the evolution of clusters within sequential data, we present a case study using posts from the AI site in the 2021 Stack Exchange dataset [41]. We apply our mini-batch k-LLMmeans algorithm with three equal-length batches and a total of ten clusters. We use the instruction “The following is a cluster of questions from the AI community. Write a single question that represents the cluster”.
Clustering results. The resulting LLM-based centroids span key areas in AI and ML, including neural network training optimization, computer vision tasks, and broader topics such as small datasets, class imbalance, and interpretability. Other clusters cover advanced themes like symbolic-neural integration for AGI, deep reinforcement learning, theoretical debates (e.g., Bayesian vs. frequentist methods), as well as NLP and various architectural choices.
Interpretation. The key insight from our algorithm is the evolution of clusters over batches, reflecting their dynamic nature over time. Figure 3 illustrates this progression for four major themes: Image Model Optimization, AI Evolution and Challenges, AI Mathematics, and Advanced NLP Techniques. The Image Model Optimization cluster refines its focus from broad improvements in image classification and object detection to specific deep learning optimizations, such as facial expression recognition and anomaly detection. Over time, it emphasizes practical challenges, including handling imbalanced datasets, model uncertainty, and performance enhancement through transfer learning and specialized loss functions. The AI Evolution and Challenges cluster transitions from early rule-based AI and symbolic systems to advanced deep learning and generative models, highlighting both progress and limitations. The shift toward hybrid and neuro-symbolic AI reflects the growing need to integrate diverse techniques for AGI, balancing efficiency, adaptability, and safety. The AI Mathematics cluster starts with unconventional symbol usage and parameter definitions in machine learning, later expanding to matrix rank analysis, Bayesian and frequentist methods, adversarial attacks, and the mathematical foundations necessary for advanced AI concepts. The Advanced NLP Techniques cluster progresses from foundational models (e.g., Word2Vec, BERT) and basic tasks (e.g., sentiment analysis, text classification) to more complex challenges, including model evaluation, domain-specific vocabulary, and multilingual alignment. The shift highlights an increasing focus on contextual understanding, transfer learning, and alternative evaluation metrics for robust NLP.
The result of this case study could enhance post categorization, searchability, answer relevance, and trend detection, making AI discussions more efficient and insightful. More broadly, this demonstrates how our interpretable mini-batch k-LLMmeans clustering algorithm, applied to sequential text streaming, can help practitioners track topic evolution, improve decision-making, and enhance transparency in dynamic information flows.
Figure 3: Sequential evolution of the LLM-generated centroids for four primary clusters during the three batches of the sequential mini-batch k-LLMmeans process applied to 2021 posts from the AI Stack Exchange site [41]. Main aspects are manually highlighted in different color on each cluster.
8 Related work
Traditional clustering. Hierarchical methods [22, 5] build tree-structured representations of nested document relationships. Density-based approaches like DBSCAN [14] and graph-based methods detect clusters of arbitrary shapes, while spectral clustering [29] leverages eigen-decomposition to uncover complex structures. Model-based techniques—including Gaussian mixture models [11] and recent neural network frameworks [52, 18, 49, 50, 48]—provide probabilistic clustering formulations. Additionally, topic modeling methods, from probabilistic latent semantic analysis [17] to latent Dirichlet allocation [6], capture word co-occurrence patterns and latent topics.
LLM-based clustering. Viswanathan et al. [45] employ LLMs to augment document representations, generate pseudo pairwise constraints, and post-correct low-confidence assignments for query-efficient, few-shot semi-supervised clustering. Zhang et al. [51] propose ClusterLLM, which uses instruction-tuned LLMs via interactive triplet and pairwise feedback to cost-effectively refine clustering granularity. Complementary approaches [44, 33] show that context-derived representations capture subtle semantic nuances beyond traditional embeddings. Wang et al. [47] introduce a goal-driven, explainable method that employs natural language descriptions to clarify cluster boundaries, while [10] present IDAS for intent discovery using abstractive summarization. Feng et al. [15] propose LLMEdgeRefine an iterative mechanism that forms super-points to mitigate outliers and reassign ambiguous edge points, resulting in clusters with higher coherence and robustness.
9 Discussion
We introduced k-LLMmeans and mini-batch k-LLMmeans, unsupervised text clustering algorithms that refine k-means by integrating a lightweight LLM-based summarization into centroid updates. This modification enriches the contextual representation while preserving k-means’ stability and scalability, yielding interpretable clusters for both static corpus of documents and streaming text data, with low LLM usage.
Limitations. Our approach depends on pre-computed embeddings and LLM-generated summaries, so any bias or error in the underlying language models can propagate to the final clusters. While we demonstrate that a simple instruction-based approach is effective, the design of these instructions can still influence outcomes. Additionally, the few-shot variant of our algorithm assumes that a small number of samples can sufficiently capture the overall structure of the clusters. While this is generally practical, it may become a limitation when dealing with highly complex or heterogeneous clusters.
Appendix A Datasets
A.1 Non-streaming datasets
We evaluate our clustering approach on four benchmark datasets. Bank77 [9]: Consists of 3,080 customer queries related to banking services, categorized into 77 distinct intents. CLINC [24]: A diverse set of 4,500 queries spanning 150 intent classes across multiple domains, designed for open-domain intent classification. GoEmo [12]: Contains 2,984 social media posts annotated with 27 fine-grained emotion categories. We removed the neutral expressions to address data imbalance and retained only entries with a single, unique emotion. MASSIVE [16]: Comprises 2,974 English-language virtual assistant utterances grouped into 18 domains and 59 intent categories.
These datasets provide a robust evaluation setting for text clustering across different domains and classification granularities.
A.2 New compiled dataset for testing text-Streaming Clustering Algorithms
We extract and unify a challenging data stream comprising unique archive posts collected from 84 Stack Exchange sites [41]. Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods. Our raw dataset spans 84 domains, each containing at least 20 posts per year from 2018 to 2023 (with post lengths ranging from 20 to 1000 characters), totaling 499,359 posts. For our experiments, we focus on posts from 2020 to 2023 and further filter out labels that do not exceed 500 posts in 2023. The resulting subset comprises 35 distinct groups and 69,147 posts. Both the raw and clean data are provided with this paper. Stack Exchange content is licensed under the Creative Commons Attribution-ShareAlike 4.0 International (CC BY-SA 4.0) license.
Appendix B Extended simulation results
B.1 Static experiments
Tables 59 present extended results from the simulations described in Section 5. For each dataset, we report the average Accuracy (ACC), Normalized Mutual Information (NMI), and distance between final and true centroids ($dist$), along with standard deviations computed over 10 random seeds. Results are provided for all k-LLMmeans variants using GPT-4o, as well as for traditional clustering algorithms, evaluated under four different embedding types. Table 10 show extended results for all benchmark datasets and multiple LLMs.
B.2 Streaming (mini-batch) experiments
Table 11 report the standard deviation for results presented on Table 4.
Appendix C Computational resources for experiments
Our simulations span approximately one week on a standard laptop without specialized hardware, with the primary computational bottleneck arising from LLM usage. We conducted API-based experiments using GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3. Most experiments relied on OpenAI’s API (GPT-4o and text-embedding-3-small), incurring a cost of roughly \$700 and consuming around 120 million tokens. Experiments involving the other LLMs collectively cost under \$30.
Appendix D Broader Impacts
While k-LLMmeans is primarily a foundational method aimed at improving scalability and interpretability in LLM-based text clustering, potential societal risks arise from the misuse of generated cluster summaries. For example, if used in sensitive domains such as legal, medical, or financial text analysis, misleading or biased summaries could propagate incorrect conclusions. Additionally, because the method relies on pretrained language models, any biases or factual inaccuracies embedded in those models may influence clustering outcomes. To mitigate these risks, we recommend cautious deployment in high-stakes settings, transparency about the LLMs used, and regular human-in-the-loop validation when interpretability affects downstream decisions.
Table 5: Average accuracy (ACC), normalized mutual information (NMI), and distance between final and true centroids (dist) for all k-LLMmeans variants with GPT-4o and all traditional clustering algorithms, evaluated on Bank77 using four different embedding models (standard deviations in parenthesis).
embedding/Method ACC NMI $dist$
DistilBERT
k-LLMmeans
   single 43.8(1.12) 64.8(0.493) 0.351(0.013)
   multiple 43.5(0.991) 64.8(0.516) 0.358(0.006)
   FS-single 43.7(0.75) 64.8(0.48) 0.352(0.007)
   FS-multiple 43.8(1.032) 64.9(0.362) 0.363(0.011)
k-means 43.653(0.714) 64.389(0.456) 0.335(0.007)
k-medoids 32.89(0.0) 55.991(0.0) 0.591(0.0)
GMM 44.701(1.163) 64.481(0.63) 0.318(0.005)
Agglomerative 45.974(0.0) 67.064(0.0) 0.33(0.0)
Spectral 45.925(0.201) 66.057(0.241) 0.365(0.001)
text-embedding-3-small
k-LLMmeans
   single 67.1(1.216) 83.6(0.459) 0.221(0.012)
   multiple 66.7(0.778) 83.6(0.2) 0.226(0.009)
   FS-single 67.5(1.072) 83.8(0.358) 0.219(0.011)
   FS-multiple 67.9(1.546) 84.1(0.412) 0.22(0.014)
k-means 66.25(1.815) 83.005(0.559) 0.225(0.014)
k-medoids 41.688(0.0) 69.457(0.0) 0.639(0.0)
GMM 68.792(1.784) 83.048(0.543) 0.198(0.013)
Agglomerative 69.87(0.0) 83.741(0.0) 0.2(0.0)
Spectral 68.562(0.562) 83.746(0.38) 0.227(0.003)
e5-large
k-LLMmeans
   single 61.1(1.041) 78.6(0.412) 0.138(0.006)
   multiple 61.6(0.783) 79.0(0.355) 0.136(0.005)
   FS-single 61.5(1.218) 78.8(0.484) 0.136(0.004)
   FS-multiple 62.7(1.706) 79.5(0.646) 0.134(0.007)
k-means 58.097(1.732) 76.703(0.795) 0.142(0.007)
k-medoids 37.175(0.0) 64.487(0.0) 0.331(0.0)
GMM 56.977(1.384) 75.801(0.502) 0.142(0.006)
Agglomerative 61.623(0.0) 79.259(0.0) 0.138(0.0)
Spectral 63.195(0.69) 79.684(0.312) 0.144(0.003)
S-BERT
k-LLMmeans
   single 64.4(1.83) 81.7(0.563) 0.253(0.014)
   multiple 64.6(1.393) 81.8(0.389) 0.253(0.013)
   FS-single 65.4(1.693) 82.0(0.515) 0.246(0.013)
   FS-multiple 65.7(1.841) 82.2(0.648) 0.247(0.013)
k-means 63.643(2.272) 80.937(0.744) 0.255(0.016)
k-medoids 41.299(0.0) 68.985(0.0) 0.64(0.0)
GMM 63.565(2.014) 80.504(0.579) 0.248(0.017)
Agglomerative 64.643(0.0) 81.601(0.0) 0.255(0.0)
Spectral 65.737(0.652) 83.247(0.395) 0.268(0.006)
Table 6: Average accuracy (ACC), normalized mutual information (NMI), and distance between final and true centroids ($dist$) for all k-LLMmeans variants with GPT-4o and all traditional clustering algorithms, evaluated on CLINC using four different embedding models (standard deviations in parenthesis).
embedding/Method ACC NMI $dist$
DistilBERT
k-LLMmeans
   single 53.3(1.468) 77.4(0.547) 0.352(0.01)
   multiple 53.5(1.141) 77.7(0.437) 0.356(0.01)
   FS-single 54.7(1.117) 78.1(0.552) 0.346(0.008)
   FS-multiple 55.3(1.104) 78.7(0.531) 0.343(0.009)
k-means 53.889(0.956) 77.188(0.494) 0.34(0.007)
k-medoids 36.889(0.0) 66.765(0.0) 0.69(0.0)
GMM 53.42(1.362) 76.534(0.442) 0.334(0.01)
Agglomerative 55.333(0.0) 78.805(0.0) 0.343(0.0)
Spectral 52.289(0.395) 76.215(0.268) 0.364(0.004)
text-embedding-3-small
k-LLMmeans
   single 78.1(1.67) 92.5(0.474) 0.194(0.015)
   multiple 79.1(1.675) 92.8(0.48) 0.186(0.015)
   FS-single 79.2(1.725) 92.8(0.427) 0.186(0.015)
   FS-multiple 80.2(1.523) 93.1(0.323) 0.179(0.015)
k-means 77.253(1.828) 91.965(0.351) 0.2(0.014)
k-medoids 49.289(0.0) 77.718(0.0) 0.694(0.0)
GMM 80.484(1.181) 92.492(0.222) 0.168(0.011)
Agglomerative 81.0(0.0) 92.544(0.0) 0.169(0.0)
Spectral 76.02(0.401) 90.812(0.139) 0.214(0.003)
e5-large
k-LLMmeans
   single 75.2(1.917) 91.6(0.48) 0.127(0.008)
   multiple 75.4(1.858) 92.0(0.472) 0.127(0.009)
   FS-single 75.4(1.746) 91.7(0.46) 0.127(0.008)
   FS-multiple 77.2(1.604) 92.5(0.399) 0.119(0.007)
k-means 73.831(2.338) 90.77(0.739) 0.131(0.01)
k-medoids 41.911(0.0) 73.371(0.0) 0.371(0.0)
GMM 77.647(1.286) 91.41(0.279) 0.106(0.006)
Agglomerative 80.333(0.0) 92.544(0.0) 0.101(0.0)
Spectral 72.853(0.252) 89.828(0.092) 0.128(0.002)
S-BERT
k-LLMmeans
   single 77.3(1.475) 91.6(0.381) 0.215(0.013)
   multiple 77.9(1.373) 91.9(0.436) 0.21(0.012)
   FS-single 78.0(1.115) 91.8(0.332) 0.208(0.01)
   FS-multiple 79.7(0.734) 92.5(0.247) 0.198(0.008)
k-means 76.853(1.63) 91.036(0.496) 0.215(0.013)
k-medoids 51.156(0.0) 78.506(0.0) 0.68(0.0)
GMM 77.333(0.93) 91.025(0.254) 0.203(0.009)
Agglomerative 82.622(0.0) 92.78(0.0) 0.158(0.0)
Spectral 77.167(0.36) 90.794(0.105) 0.215(0.002)
Table 7: Average accuracy (ACC), normalized mutual information (NMI), and distance between final and true centroids ($dist$) for all k-LLMmeans variants with GPT-4o and all traditional clustering algorithms, evaluated on GoEmo using four different embedding models (standard deviations in parenthesis).
embedding/Method ACC NMI $dist$
DistilBERT
k-LLMmeans
   single 17.5(0.467) 18.3(0.447) 0.358(0.009)
   multiple 17.9(0.861) 18.3(0.62) 0.354(0.006)
   FS-single 18.0(0.681) 18.9(0.407) 0.357(0.006)
   FS-multiple 18.2(0.52) 18.8(0.733) 0.351(0.004)
k-means 17.755(0.857) 18.267(0.667) 0.364(0.005)
k-medoids 15.851(0.0) 13.964(0.0) 0.769(0.0)
GMM 18.07(0.819) 18.729(0.471) 0.36(0.005)
Agglomerative 16.857(0.0) 15.402(0.0) 0.352(0.0)
Spectral 17.198(0.21) 17.23(0.189) 0.42(0.002)
text-embedding-3-small
k-LLMmeans
   single 24.0(1.346) 22.3(0.572) 0.278(0.007)
   multiple 24.1(1.326) 22.1(0.651) 0.275(0.004)
   FS-single 23.0(0.937) 21.9(0.638) 0.283(0.011)
   FS-multiple 24.2(1.156) 22.3(0.67) 0.278(0.011)
k-means 20.734(0.997) 20.466(0.672) 0.287(0.005)
k-medoids 15.65(0.0) 15.943(0.0) 0.852(0.0)
GMM 20.831(1.418) 20.123(0.524) 0.296(0.007)
Agglomerative 15.818(0.0) 14.098(0.0) 0.286(0.0)
Spectral 17.661(0.328) 15.045(0.319) 0.341(0.004)
e5-large
k-LLMmeans
   single 23.4(0.875) 23.4(0.694) 0.176(0.007)
   multiple 23.4(0.9) 23.7(0.521) 0.174(0.004)
   FS-single 23.4(1.089) 23.7(0.356) 0.174(0.003)
   FS-multiple 24.2(0.761) 24.3(0.702) 0.168(0.003)
k-means 22.778(0.917) 22.761(0.729) 0.176(0.003)
k-medoids 14.812(0.0) 15.667(0.0) 0.447(0.0)
GMM 21.367(0.986) 22.197(0.69) 0.181(0.003)
Agglomerative 19.571(0.0) 18.544(0.0) 0.177(0.0)
Spectral 23.388(0.524) 21.362(0.257) 0.192(0.001)
S-BERT
k-LLMmeans
   single 14.4(0.739) 13.6(0.547) 0.351(0.006)
   multiple 14.4(0.75) 13.7(0.59) 0.348(0.005)
   FS-single 14.5(0.79) 13.9(0.413) 0.348(0.006)
   FS-multiple 14.7(0.551) 13.9(0.501) 0.346(0.004)
k-means 13.737(0.617) 13.269(0.574) 0.355(0.005)
k-medoids 13.941(0.0) 11.974(0.0) 0.899(0.0)
GMM 14.31(0.744) 13.628(0.482) 0.353(0.003)
Agglomerative 12.399(0.0) 10.766(0.0) 0.355(0.0)
Spectral 13.05(0.211) 11.777(0.123) 0.416(0.003)
Table 8: Average accuracy (ACC), normalized mutual information (NMI), and distance between final and true centroids ($dist$) for all k-LLMmeans variants with GPT-4o and all traditional clustering algorithms, evaluated on MASSIVE (Domain) using four different embedding models (standard deviations in parenthesis).
embedding/Method ACC NMI $dist$
DistilBERT
k-LLMmeans
   single 45.8(1.508) 46.0(1.164) 0.294(0.016)
   multiple 45.2(2.264) 45.6(1.199) 0.306(0.017)
   FS-single 45.3(1.789) 45.7(1.099) 0.299(0.021)
   FS-multiple 46.2(1.348) 46.4(1.119) 0.295(0.015)
k-means 44.368(1.528) 45.065(0.988) 0.309(0.021)
k-medoids 33.457(0.0) 27.223(0.0) 0.81(0.0)
GMM 45.451(1.892) 46.966(0.74) 0.293(0.02)
Agglomerative 35.777(0.0) 39.931(0.0) 0.355(0.0)
Spectral 40.212(2.693) 44.476(1.514) 0.473(0.009)
text-embedding-3-small
k-LLMmeans
   single 61.1(3.393) 69.6(1.604) 0.232(0.018)
   multiple 60.6(2.822) 69.5(1.661) 0.238(0.016)
   FS-single 60.8(3.206) 69.4(1.091) 0.233(0.02)
   FS-multiple 63.2(2.843) 70.6(1.379) 0.227(0.022)
k-means 59.391(4.026) 67.869(1.671) 0.246(0.015)
k-medoids 37.559(0.0) 38.788(0.0) 0.797(0.0)
GMM 59.391(2.883) 68.0(1.01) 0.253(0.014)
Agglomerative 62.845(0.0) 67.07(0.0) 0.222(0.0)
Spectral 61.51(0.022) 66.987(0.023) 0.331(0.0)
e5-large
k-LLMmeans
   single 60.6(2.856) 64.8(1.611) 0.134(0.014)
   multiple 61.5(4.447) 66.0(2.105) 0.131(0.016)
   FS-single 61.0(3.441) 64.9(1.42) 0.136(0.019)
   FS-multiple 62.3(1.358) 65.9(0.776) 0.133(0.007)
k-means 58.426(3.889) 63.663(1.748) 0.138(0.02)
k-medoids 41.19(0.0) 36.343(0.0) 0.405(0.0)
GMM 63.073(3.473) 66.018(1.588) 0.132(0.008)
Agglomerative 66.073(0.0) 66.279(0.0) 0.132(0.0)
Spectral 57.001(0.02) 63.327(0.018) 0.202(0.0)
S-BERT
k-LLMmeans
   single 61.0(3.797) 65.6(1.36) 0.248(0.026)
   multiple 60.1(2.016) 65.2(0.735) 0.25(0.019)
   FS-single 61.7(2.529) 66.2(1.134) 0.243(0.026)
   FS-multiple 59.7(2.543) 65.6(0.795) 0.254(0.019)
k-means 58.248(2.029) 64.608(1.09) 0.271(0.023)
k-medoids 45.763(0.0) 44.605(0.0) 0.8(0.0)
GMM 57.935(3.645) 64.404(1.094) 0.27(0.032)
Agglomerative 63.416(0.0) 64.72(0.0) 0.222(0.0)
Spectral 58.487(0.85) 63.148(0.49) 0.367(0.005)
Table 9: Average accuracy (ACC), normalized mutual information (NMI), and distance between final and true centroids ($dist$) for all k-LLMmeans variants with GPT-4o and all traditional clustering algorithms, evaluated on MASSIVE (Intention) using four different embedding models (standard deviations in parenthesis).
embedding/Method ACC NMI $dist$
DistilBERT
k-LLMmeans
   single 44.1(1.099) 58.6(0.511) 0.362(0.014)
   multiple 43.8(1.494) 59.0(0.46) 0.363(0.013)
   FS-single 43.5(1.196) 58.8(0.59) 0.365(0.01)
   FS-multiple 43.6(1.343) 59.0(0.594) 0.362(0.012)
k-means 43.124(1.465) 58.102(0.516) 0.366(0.014)
k-medoids 25.084(0.0) 42.582(0.0) 0.847(0.0)
GMM 41.379(1.249) 57.375(0.452) 0.367(0.013)
Agglomerative 39.139(0.0) 55.378(0.0) 0.386(0.0)
Spectral 46.022(0.482) 58.237(0.234) 0.378(0.004)
text-embedding-3-small
k-LLMmeans
   single 54.0(1.513) 73.0(0.582) 0.314(0.013)
   multiple 55.8(1.396) 73.5(0.762) 0.307(0.013)
   FS-single 55.3(1.759) 73.6(0.774) 0.314(0.014)
   FS-multiple 56.4(1.32) 73.9(0.7) 0.302(0.012)
k-means 52.878(1.072) 72.373(0.541) 0.32(0.011)
k-medoids 35.911(0.0) 52.442(0.0) 0.86(0.0)
GMM 54.56(1.918) 73.105(0.798) 0.306(0.015)
Agglomerative 56.59(0.0) 70.461(0.0) 0.29(0.0)
Spectral 58.453(0.837) 71.797(0.23) 0.28(0.002)
e5-large
k-LLMmeans
   single 55.2(1.625) 72.0(0.515) 0.17(0.01)
   multiple 55.6(2.161) 72.1(0.725) 0.166(0.012)
   FS-single 54.3(2.395) 71.6(0.833) 0.173(0.012)
   FS-multiple 55.8(1.995) 72.4(0.912) 0.166(0.01)
k-means 52.885(1.264) 70.917(0.708) 0.175(0.01)
k-medoids 32.112(0.0) 50.313(0.0) 0.437(0.0)
GMM 50.595(2.28) 70.15(1.225) 0.183(0.007)
Agglomerative 55.447(0.0) 70.315(0.0) 0.17(0.0)
Spectral 56.594(0.585) 69.034(0.417) 0.152(0.003)
S-BERT
k-LLMmeans
   single 52.2(2.754) 71.2(0.785) 0.334(0.022)
   multiple 52.9(2.005) 71.5(0.675) 0.325(0.015)
   FS-single 52.0(2.552) 71.4(0.507) 0.333(0.019)
   FS-multiple 53.3(1.768) 71.9(0.466) 0.337(0.011)
k-means 50.908(1.747) 70.708(0.296) 0.344(0.019)
k-medoids 35.978(0.0) 54.587(0.0) 0.817(0.0)
GMM 50.494(1.22) 70.559(0.488) 0.338(0.012)
Agglomerative 52.186(0.0) 69.544(0.0) 0.313(0.0)
Spectral 56.126(1.395) 70.227(0.463) 0.304(0.005)
Table 10: Average accuracy (ACC) and normalized mutual information (NMI) for k-LLMmeans FS-multiple using various LLMs and e5-large embeddings on all benchmark datasets (standard deviations in parenthesis).
Dataset/Method Bank77 CLINC GoEmo Massive (D) Massive (I)
ACC NMI ACC NMI ACC NMI ACC NMI ACC NMI
k-LLMmeans
   GPT-3.5 56.6 77.6 76.0 92.1 23.9 24.1 63.3 66.2 56.9 72.5
(1.32) (0.67) (1.16) (0.28) (0.65) (0.39) (1.47) (0.62) (1.62) (0.86)
   GPT-4o 62.7 79.5 77.2 92.5 24.2 24.3 62.3 65.9 55.8 72.4
(1.71) (0.65) (1.6) (0.4) (0.76) (0.7) (1.36) (0.78) (1.99) (0.91)
   Llama-3.3 61.6 79.2 77.2 92.3 23.8 23.1 62.0 66.3 56.8 72.9
(1.7) (0.63) (1.6) (0.44) (0.66) (0.62) (2.97) (1.56) (1.26) (0.5)
   DeepSeek-V3 63.1 79.9 69.7 90.8 22.8 22.7 62.6 66.3 57.6 72.0
(1.61) (0.75) (1.02) (0.29) (1.68) (0.88) (2.24) (0.71) (1.43) (0.61)
   Claude-3.7 61.0 78.8 76.9 92.5 24.2 23.7 61.8 66.0 57.3 72.9
(1.52) (0.53) (1.83) (0.52) (1.0) (0.65) (4.83) (1.99) (1.47) (0.85)
Table 11: Standard deviation of the accuracy (ACC) and Normalized Mutual Information (NMI) for four sequential mini-batch variants, k-means, mini-batch k-means, sequential mini-batch k-means on the yearly StackExchange data.
Year/Method 2020 2021 2022 2023
(69147 posts) (54322 posts) (43521 posts) (38953 posts)
ACC NMI ACC NMI ACC NMI ACC NMI
mini-batch k-LLMmeans
   single (3.19) (0.57) (2.48) (0.42) (2.72) (0.86) (1.59) (0.46)
   multiple (3.02) (0.65) (2.18) (0.46) (2.48) (0.87) (2.24) (0.81)
   FS + single (2.86) (0.39) (2.15) (0.35) (2.68) (0.9) (1.68) (0.59)
   FS + multiple (2.02) (0.48) (2.39) (0.7) (2.28) (0.91) (1.38) (0.45)
k-means (3.79) (0.75) (2.8) (0.25) (4.03) (0.69) (3.79) (0.56)
mini-batch k-means (2.09) (0.9) (2.12) (0.84) (4.02) (0.98) (2.76) (0.97)
seq. mini-batch k-means (2.39) (0.57) (4.61) (1.54) (1.23) (0.78) (2.55) (0.9)